# Lossy Compression with Data, Perception, and Classification Constraints

### v1

@@ Abstract
Balancing diverse task objectives under limited rate is crucial for developing robust multi-task deep learning (DL) models and improving performance across various domains. In this paper, we consider the lossy compression problem with human-centric and task-oriented metrics, such as perceptual quality and classification accuracy. We investigate two ternary relationships, namely, the rate-distortion-classification (RDC) and rate-perception-classification (RPC). For both RDC and RPC functions, we derive the closed-form expressions of the optimal rate for both binary and Gaussian sources. Notably, both RDC and RPC relationships exhibit distinct characteristics compared to the previous RDP tradeoff proposed by Blau et al. Then, we conduct experiments by implementing a DL-based image compression framework, incorporating rate, distortion, perception, and classification constraints. The experimental results verify the theoretical characteristics of RDC and RPC tradeoffs, providing information-theoretical insights into the design of loss functions to balance diverse task objectives in deep learning.
@@ I Introduction
Towards an information-theoretical explanation of deep learning (DL) algorithms, the information bottleneck (IB) principle[1] has received significant attention. The IB principle introduces an additional variable that determines relevance, aiming to find a maximally compressed mapping of the input variable that preserves relevant information. This principle is connected to a remote lossy source-coding problem with logarithmic loss distortion[2, 3]. Since its introduction in 1999, the IB principle has been adapted to a broad spectrum of application domains in machine learning[4], offering a novel paradigm for analyzing deep neural networks on their layered structure[4, 5]. Meanwhile, the idea of IB has been successfully introduced to the design of communication-efficient edge inference systems[6, 7].
While the IB principle primarily focuses on tasks of target label inference, recent research has also introduced the concept of perception quality to rate-distortion analysis, thereby enabling an information-theoretical understanding of generative tasks. In the study conducted by [8], a mathematical proxy to measure perception quality based on the divergence between the source and reconstruction distributions is employed. The authors proposed the rate-distortion-perception (RDP) function, analyzing the ternary tradeoff between rate, distortion, and perception. The follow-up work [9] derived the closed-form expression of RDP function for the Gaussian source under squared error distortion and Wasserstein-2 perception loss. Meanwhile, different coding schemes have been established[10, 11, 12], showing that the RDP function can be achieved using stochastic or deterministic codes.
The tradeoffs between different tasks have been explored from the perspective of signal degradation and restoration. These studies typically involve signals that are corrupted by extrinsic noise, and only a degraded version of the source signal is available for reconstruction or denoising. In the study of perception-distortion (DP) tradeoff [13], it is demonstrated that improvements in distortion do not necessarily align with improvements in visual quality. Furthermore, in [14], classification accuracy is considered along with perceptual quality and distortion. They considered the classification error rate as a representative of semantic quality and studied classification-distortion-perception (CDP) tradeoff. The experiments showed that better classification performance comes at the cost of higher distortion or worse perception.
In this paper, we extend the scope of the aforementioned research and obtain a deeper understanding of the interplay between rate, distortion, perception, and classification. Specifically, we address two open problems: the information rate-distortion-classification (RDC) function and the information rate-perception-classification (RPC) function. For both RDC and RPC functions, we derive the closed-form expressions for the optimal rate in both binary and Gaussian scenarios. Interestingly, we find that RDC, RPC, and the previous RDP[8] exhibit distinct characteristics in each trinary relationship. We also verify our theoretical results by implementing a DL-based image compression framework that incorporates diverse constraints. As new results in characterizing task-relevant information, RDC and RPC functions complement the IB theory and RDP tradeoff, and provide information-theoretical insights into designing loss functions to balance diverse task objectives in deep learning.
Table I provides a summary of the tradeoffs between distortion, perception, and classification in both lossy compression and signal restoration frameworks, as presented in previous literature and this paper.
@@
Reference Tradeoff Closed form
[15] RD Binary[15]
@@
Mixed Gaussian[14]
(Simulation only)
@@
Note: Here R, D, P, and C refers to rate, distortion,
perception, and classification respectively.
@@ II Problem Formulation
Consider a source with observable data $X\sim p_{X}(x)$, which intrinsically contains several classification aspects formulated by variables $S_{1},\cdots,S_{K}\sim p_{S}(s_{1},\cdots,s_{K})$. The observation $X$ and the intrinsic variables are correlated and follow the joint probability distribution $p_{X,S}(x,s_{1},\cdots,s_{K})$ over $\mathcal{X}\times\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{K}$. The intrinsic variable is not observable but can be inferred from $X$. For example, the observation $X$ could be a voice signal, and the interested classification variables may be the transcription of the speech, the identity of the speaker, or the gender of the speaker.
@@
As shown in the Fig. 1, the process of lossy compression consists of an encoder and a decoder.
  • The encoding function $f:\mathcal{X}\mapsto\mathcal{Z}$ maps the observable source $X$ to an efficient representation $Z=f(X)$.
  • The decoding function $g:\mathcal{Z}\mapsto\mathcal{\hat{X}}$ reproduces the source data $\hat{X}$ to satisfy task-oriented demands of downstream applications.
The joint distribution of $X$ and $\hat{X}$ follows $p_{X,\hat{X}}(x,\hat{x})=p_{\hat{X}|X}(\hat{x}|x)p_{X}(x)$. Meanwhile, $(S_{1},\cdots,S_{K})$, $X$ and $\hat{X}$ follow the Markov chain $(S_{1},\cdots,S_{K})\rightarrow X\rightarrow\hat{X}$.
In task-oriented lossy compression, the destination may involve various tasks upon receiving a compressed signal. These tasks can include traditional data reconstruction, generative learning tasks, or the extraction of classification labels. As a result, the following constraints arise in order to accommodate these potential applications:
1) Reconstruction constraint: We consider the following data reconstruction constraint
$\mathbb{E}[\Delta(X,\hat{X})]\leq D,$
where $\Delta:\mathcal{X}\times\hat{\mathcal{X}}\to\mathbb{R}^{+}$ is a distortion function, such as Hamming distortion and squared-error distortion, and the expectation is taken over $p_{X,\hat{X}}(x,\hat{x})=p_{\hat{X}|X}(\hat{x}|x)p_{X}(x)$. In the subsequent discussion, the term "distortion" specifically refers to the data reconstruction distortion described here.
2) Perception constraint: The perceptual quality usually refers to the degree to which an image resembles a natural image rather than a synthetic or restored image generated by an algorithm[18]. It has been demonstrated that perceptual quality is associated with the distance between the distributions of natural images and generated images [13]. Meanwhile, the underlying principle of generative adversarial network (GAN)[19] and its variants (e.g., Wasserstein GAN[20]) also involves minimizing a divergence between two distributions. Hence, in this paper, we adopt the same perception constraint as [13, 8, 9, 14], i.e.,
@@
3) Classification constraint: The conditional entropy $H(X|Y)$ measures the uncertainty in $X$ given information $Y$, which also serves as a lower bound for the cross-entropy loss[21], a commonly employed metric in machine learning for optimizing classification models. Thus, we adopt the following classification constraint
@@
for some $C_{k}>0$, which means the uncertainty of classification variable $S_{k}$ given the recovered source $\hat{X}$ should not surpass the level $C_{k}$. Equivalently, the constraints can be written as $I(S_{k},\hat{X})\geq C^{{}^{\prime}}_{k},\forall k\in[K],$ which indicates that a certain amount of semantic information about the relevant variable $S_{k}$ is preserved in $\hat{X}$.
Incorporating distortion, perception, and classification constraints, we define the information rate-distortion-perception-classification (RDPC) function for a source $X\sim p_{X}(x)$ as
@@
Connection with previous work: The proposed $R(D,P,\mathbf{C})$ function can be reduced to a series of previous theories by relaxing one or two constraints: When the classification constraint (1c) is relaxed, $R(D,P,\mathbf{C})$ is equivalent to the RDP function in [8]. If the perception constraint is also relaxed, the RDP function is reduced to the rate-distortion function $R(D)$ [15, 22]. When there exists only a single classification variable and both reconstruction (1a) and perception constraints (1b) are inactive, the proposed $R(D,P,\mathbf{C})$ is reduced to
$\min_{p_{\hat{X}|X}:I(S,\hat{X})\geq C}I(X;\hat{X})$
By introducing the Lagrange multiplier $\beta$, the above problem is equivalent to $\min_{p_{\hat{X}|X}}\{I(X;\hat{X})-\beta I(S,\hat{X})\}$, which is the information bottleneck problem[1].
In the following sections, we will delve into a theoretical analysis of two trinary relationships: rate-distortion classification (RDC) and rate-perception-classification (RPC), which have not been investigated by previous literature. We will provide closed-form expressions for both binary and Gaussian sources, and explore distinctions between different tradeoffs.
III Rate-distortion-classification Tradeoff
In many practical scenarios, reconstruction distortion is concerned jointly with classification accuracies, such as autonomous driving and face recognition in surveillance applications[23, 24]. In this section, we characterize the inherent RDC tradeoff and derive closed-form expression of the optimal rate for binary and Gaussian sources.
When perception constraint (1b) is relaxed and consider only one classification variable, we can obtain the following information rate-distortion-classification (RDC) function:
@@ III Rate-distortion-classification Tradeoff
III-A Binary Source
For the binary source, we characterize the closed-form solution for $R(D,C)$ as the following theorem.
@@ III-A Binary Source
For a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$), if $C<H(S_{1})$, the problem is infeasible; otherwise, the information rate-distortion-classification function with Hamming distortion is given by
@@
From the theorem, we can observe that when the distortion constraint $D$ is relatively large, the rate $R$ becomes a function of $C$, indicating that the classification constraint becomes the primary limiting factor. Conversely, when $D$ is relatively small and becomes the main limiting factor, the rate $R$ becomes a function of the distortion constraint $D$.
In Fig. 2(a), we visualize the result of Theorem 1 by plotting the $R(D,C)$ function of a Bernoulli source as a surface (Fig. 2(a)) as well as distortion-classification curves for different rates (Fig. 2(b)). The first observation from the figures is that $R(D,C)$ function is non-increasing and convex over $D$ and $C$. When $D$ is relatively large, the value of rate is primarily determined by the value of $C$, and vice versa. When rate bits become larger, the overall curve shifts to the lower left, which means we can achieve a better classification performance and a better distortion level simultaneously.
III-B Gaussian Source

### v2

@@ Abstract
By extracting task-relevant information while maximally compressing the input, the information bottleneck (IB) principle has provided a guideline for learning effective and robust representations of the target inference. However, extending the idea to the multi-task learning scenario with joint consideration of generative tasks and traditional reconstruction tasks remains unexplored. This paper addresses this gap by reconsidering the lossy compression problem with diverse constraints on data reconstruction, perceptual quality, and classification accuracy. Firstly, we study two ternary relationships, namely, the rate-distortion-classification (RDC) and rate-perception-classification (RPC). For both RDC and RPC functions, we derive the closed-form expressions of the optimal rate for binary and Gaussian sources. These new results complement the IB principle and provide insights into effectively extracting task-oriented information to fulfill diverse objectives. Secondly, unlike prior research demonstrating a tradeoff between classification and perception in signal restoration problems, we prove that such a tradeoff does not exist in the RPC function and reveal that the source noise plays a decisive role in the classification-perception tradeoff. Finally, we implement a deep-learning-based image compression framework, incorporating multiple tasks related to distortion, perception, and classification. The experimental results coincide with the theoretical analysis and verify the effectiveness of our generalized IB in balancing various task objectives.
Index Terms:
Information bottleneck, lossy compression, task-oriented communication, rate-distortion theory, perceptual quality.
@@ I Introduction
The last decades have witnessed significant achievements in machine learning, particularly with the success of deep learning (DL) methods across various tasks. Leveraging information-theoretical insights can facilitate the understanding of DL algorithms and aid in designing loss functions. The information bottleneck (IB) principle[1] emerged as a promising theory for analyzing and training DL algorithms, especially in the context of feature extraction. The IB principle aims to find a compressed mapping $\hat{X}$ of the input observation $X$ while preserving enough information about a correlated target $S$. Specifically, the IB problem is given by
$\min_{p_{\hat{X}|X}:I(S;\hat{X})\geq C}I(X;\hat{X}),$
where $C$ is a constant and $S-X-\hat{X}$ forms Markov chain. The IB principle can be viewed as a remote lossy source-coding problem with logarithmic loss distortion[2, 3]. Since its introduction in 1999, the IB principle has received significant attention and has been widely applied to design machine learning (e.g., representation learning) [4, 5] and task-oriented communications (e.g., edge inference) [6, 7].
On the other hand, recent research has introduced the concept of perception quality to the lossy compression problem. The authors in [8] proposed the rate-distortion-perception (RDP) function, and analyzed this ternary tradeoff. Here the perception is defined to be the divergence between the source and reconstruction distributions. In [9], the perception measure based on the divergence between distributions conditioned on the encoder output is also studied, which empirically results in higher perceptual quality of reconstructions[10]. In [11], the closed-form expression of the RDP function for the Gaussian source is derived under mean-square error (MSE) distortion and squared Wasserstein-2 distance. Notably, when considering perfect realism where the perception loss is zero, the RDP function aligns with the theory of distribution-preserving rate-distortion function studied in [12]. The cost of perfect realism in lossy compression was explored in [13], and the advantage of stochastic encoders in the one-shot setting was illustrated in [14]. Additionally, different coding schemes for the RDP tradeoff have been established recently [15, 16], demonstrating that the RDP function can be achieved by stochastic or deterministic codes.
@@
Ref. Tradeoff Closed form Note
[17] RD Binary[17] Source noise in signal restoration makes the perfect reconstruction impossible, leading to the persistence of tradeoffs even in the absence of a rate constraint. The details will be discussed in Section IV.
@@
Gaussian mixture
(Simulation)[21]
@@
RPC
given D
Gaussian
Note: R, D, P, and C refer to rate, data
distortion, perception and classification.
Similar tradeoffs have also been explored in the problem of signal degradation and restoration, where signals are corrupted by extrinsic noise and only a degraded version of the source signal is available for reconstruction or denoising. According to the study in [20][22], there is not always a direct correlation between the technical reduction of distortion and the enhancement of perceived visual quality. The authors mathematically proved the existence of a tradeoff between distortion and perception, known as the distortion-perception (DP) tradeoff. Furthermore, the consideration of classification tasks alongside perceptual quality and distortion was addressed in [21]. The authors investigated the tradeoff between classification, distortion, and perception, referred to as the classification-distortion-perception (CDP) tradeoff. The experimental results highlight that achieving better classification performance often comes at the expense of higher distortion or poorer perceptual quality.
I-A Motivations
Although existing works provide certain explanations and design guidance for DL in various tasks, theoretical analysis is still missing for many important applications. For example, reconstruction distortion is concerned jointly with classification accuracies in scenarios such as autonomous driving and face recognition in surveillance systems[23, 24]. Meanwhile, generative tasks and classification tasks are often jointly considered in the context of conditional generative adversarial nets (C-GAN)[25], InfoGAN[26], and CatGAN[27] where the loss function is usually a composition of GAN-loss and mutual information regularization. In these scenarios, there is a lack of sufficient theoretical characterization of the relationships between target inference tasks and generative tasks as well as traditional reconstruction tasks.
Furthermore, multi-task machine learning is more complex compared to single-task learning. In particular, different tasks may exhibit conflicting needs, making it crucial to explore the tradeoffs among them under limited resources[28]. Meanwhile, although several heuristic designs for multi-task loss functions have been proposed[28], no guidelines on balancing different tasks have been provided from an information-theoretical perspective. The existing IB principle has been successfully applied to single-task scenarios with a focus on feature extraction. A multi-task information bottleneck theory with joint consideration of generative tasks and even traditional restoration tasks is more desirable. Such a theory has great potential to provide insights into understanding tradeoffs in diverse tasks and offer design guidelines for multi-task learning.
I-B Our Contributions
To address the above issues, we integrate traditional reconstruction tasks and generative tasks into the IB principle and investigate both the theoretical and practical roles of the generalized IB framework. Our contributions are summarized as follows.
  • First, we investigate two open problems: the rate-distortion-classification (RDC) function and the rate-perception-classification (RPC) function. We derive the closed-form expressions for both RDC and RPC functions in binary and scalar Gaussian scenarios. As new results in characterizing task-relevant information, RDC and RPC functions complement the IB theory and RDP tradeoff, providing insights on balancing multiple tasks under a limited compression rate.
  • In contrast to existing work showing the existence of a tradeoff between classification and perception [21] in the signal degradation and restoration problem, our theoretical analysis shows that such a tradeoff does not exist in the RPC function. To account for this contradiction, we reveal that the source noise plays a decisive role, and show that once given a certain level of distortion, the classification-perception appears in the RPC function.
  • Finally, we conduct a series of experiments by implementing a DL-based image compression framework incorporating multiple tasks. The experimental outcomes validate our theoretical results on the RDC, RPC tradeoffs, as well as RPC with certain levels of distortion. The experiments demonstrate that the generalized IB framework could effectively provide guidelines for designing DL learning methods with multiple task objectives.
Table I provides a summary of the tradeoffs between distortion, perception, and classification tasks in both lossy compression and signal restoration, as presented in previous literature and this paper. The tradeoff between rate and classification (RC) refers to the IB principle[1]. Note that although the CDP tradeoff has been previously studied in [21] in the view of signal restoration, our settings differ greatly in two aspects: First, the tradeoffs between different tasks are evaluated in compressed scenarios, which could affect the behavior of all three tasks. Second, the analysis of classification performance in [21] is characterized by a predefined classifier while we use conditional entropy with broader applications in feature extraction and target inference.
The rest of this paper is organized as follows. In Section II we present the problem formulation and the proposed information rate-distortion-perception-classification function. The RDC and RPC functions are investigated in Section III. In Section IV, we discuss the decisive role of source noise in the existence of tradeoffs. Then, experimental results are presented in Section V. Finally, Section VI concludes the paper.
Notations: For a random variable $X$ denoted as a capital letter, we use small letter $x$ for its realizations, and use $p_{X}(x)$ to denote the distribution over its alphabet $\mathcal{X}$. The expectation is denoted as $\mathbb{E}[X]$. We use $H(\cdot)$ to denote the Shannon entropy of a discrete random variable, and $h(\cdot)$ to represent the differential entropy.
@@ II Problem Formulation
Consider a source with observable data $X\sim p_{X}(x)$, which intrinsically contains several target labels formulated by variables $S_{1},\cdots,S_{K}\sim p_{S}(s_{1},\cdots,s_{K})$. The observation $X$ and the intrinsic variables are correlated and follow a joint probability distribution $p_{X,S}(x,s_{1},\cdots,s_{K})$ over $\mathcal{X}\times\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{K}$. The label variables are not observable but could be inferred from $X$. For example, the observation $X$ could be a voice signal, and the interested classification variables may be the transcription of the speech, the identity of the speaker, or the gender of the speaker.
@@
As shown in the Fig. 1, the process of lossy compression consists of an encoder and a decoder. Assume we have a source producing an i.i.d. sequence $X_{1},X_{2},\cdots,X_{n}\sim p(x)$.
  • The encoding function $f:\mathcal{X}^{n}\mapsto\{1,2,\cdots,2^{nR}\}$ maps the source $X^{n}$ to a message $M$ with a rate of $R$ bits.
  • The decoding function $g:\{1,2,\cdots,2^{nR}\}\mapsto\mathcal{\hat{X}}^{n}$ reproduces data $\hat{X}^{n}$ to satisfy task-oriented demands of downstream applications.
In task-oriented lossy compression, the destination may involve various tasks upon receiving a compressed signal, including traditional data reconstruction, generative learning tasks, or the prediction of classification labels. To accommodate these potential applications, we present the following symbol-level constraints. Theoretical analysis based on these constraints could reveal the practical usefulness of the generalized IB as an optimization objective[4].
1) Reconstruction constraint: We consider the following reconstruction constraint
$\mathbb{E}(\Delta(X,\hat{X}))\leq D,$
where $\Delta:\mathcal{X}\times\hat{\mathcal{X}}\to\mathbb{R}^{+}$ is a data distortion function, such as Hamming distortion and squared-error distortion, and the expectation is taken over $p_{X,\hat{X}}(x,\hat{x})=p_{\hat{X}|X}(\hat{x}|x)p_{X}(x)$.
2) Perception constraint: The perceptual quality usually refer to the degree to which an image resembles a natural image rather than a synthetic or restored image generated by an algorithm[29]. It has been demonstrated that perceptual quality could be associated with the distance between the distributions of natural images and generated images [20]. Meanwhile, the underlying principle of generative adversarial network (GAN)[30] and its variants (e.g., Wasserstein GAN[31]) involve minimizing a divergence between two distributions. Hence, in this paper we adopt the same perception constraint as [20, 8, 11, 21]
@@
3) Classification constraint: The conditional entropy $H(X|Y)$ measures the uncertainty in $X$ given information $Y$. Here we adopt the following classification constraint
@@
for some $C_{k}>0$, which means the uncertainty of classification variable $S_{k}$ given the recovered source $\hat{X}$ should not surpass the level $C_{k}$. Equivalently, the constraints can be written as $I(S_{k},\hat{X})\geq C^{{}^{\prime}}_{k},\forall k\in[K]$, which indicates that a certain amount of semantic information about the relevant variable $S_{k}$ is preserved in $\hat{X}$. In Section V-A, we will show that conditional entropy serves as a lower bound for the cross-entropy loss, a commonly employed metric in machine learning for optimizing classification models.
To characterize the achievable rate under all distortion, perception, and classification constraints, we can define the information rate-distortion-perception-classification (RDPC) function for a source $X\sim p_{X}(x)$ as
@@
Connection with previous work: The proposed $R(D,P,\mathbf{C})$ function can be reduced to a series of previous theories by relaxing one or two constraints: When the classification constraint (1c) is relaxed, $R(D,P,\mathbf{C})$ is equivalent to the RDP function in [8]. If the perception constraint is also relaxed, the RDP function is reduced to the rate-distortion function $R(D)$ [17, 32]. When there exists only a single classification variable and both reconstruction (1a) and perception constraints (1b) are inactive, the proposed $R(D,P,\mathbf{C})$ is equivalent to information bottleneck principle[1]. Specifically, the problem is reduced to $\min_{p_{\hat{X}|X}:I(S,\hat{X})\geq C}I(X;\hat{X})$. By introducing the Lagrange multiplier $\beta$, the above problem is equivalent to $\min_{p_{\hat{X}|X}}\{I(X;\hat{X})-\beta I(S,\hat{X})\}$, which is the information bottleneck problem[1].
Remark 1 (Optimality with strong asymptotical constraints).
The operational meaning of the proposed information RDPC function can be obtained from the results in [15] via the Poisson functional representation[33] and common randomness exists between the encoder and decoder. The proof of both converse and achievability follows from demonstrating that the constraints in (1a)-(1c) are functions of the joint distribution $p_{\hat{X},X}$, thereby making the information RDPC function a specific case of the information rate function defined in [15]. Note that asymptotical achievability in [15] is very strong in the sense that all constraints should be satisfied by each single message.
Remark 2 (Achievability with weak asymptotical constraints).
In a weak definition of achievability where the constraints are satisfied averagely (e.g., $\sum_{i}^{N}\mathbb{E}[\Delta(X_{n},\hat{X}_{n})]\leq D$ for some i.i.d process $\{X_{n}\}_{1}^{N}$ with margin $p_{X}$), the RDPC function is still achievable since the strong achievability implies the weak achievability. Meanwhile, we can establish the optimality of RDC function by proving its convexity in general (See Appendix E). For the RPC function, as demonstrated in Section III-B, it reduces to the RC case for the Bernoulli and scalar Gaussian sources, which are also convex. See
Nevertheless, this paper does not primarily focus on the operational definition of the RDPC function. Instead, its main objective is to provide insights into the tradeoffs between the rate and different tasks, as well as offer guidance for balancing diverse objectives in DL algorithms through the design of loss functions. The effectiveness of the RDPC function will be demonstrated with our experimental results in Section V.
III Investigating RDC and RPC Tradeoffs
In this section, we will initially examine two unexplored relationships by relaxing one constraint of the RDPC function: the rate-distortion-classification (RDC) and the rate-perception-classification (RPC).
III-A Rate-distortion-classification Tradeoff
When relaxing perception constraint (1b) and considering only one classification variable, we obtain the following information rate-distortion-classification function:
@@ III-A Rate-distortion-classification Tradeoff
III-A1 Binary source
For binary sources, we characterize the closed-form solution for $R(D,C)$ as the following theorem.
@@ III-A1 Binary source
Consider a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). The problem is infeasible if $C<H(S_{1})$. Otherwise, the information rate-distortion-classification function with Hamming distortion is given by
@@
From the theorem, we can observe that when the distortion constraint $D$ is relatively large, the rate $R$ becomes a function of the classification constraint $C$, indicating that the classification constraint becomes the primary limiting factor. Conversely, when $D$ is relatively small and distortion becomes the dominant constraint, the rate $R$ becomes a function of the distortion constraint $D$.
In Fig. 2(a), we visualize the result of Theorem 1 by plotting the $R(D,C)$ function of a Bernoulli source as a surface (Fig. 2(a), left) as well as distortion-classification curves for different rates (Fig. 2(a), right). The first observation from the figures is that $R(D,C)$ function is non-increasing and convex over $D$ and $C$. When D is relatively large, the value of rate is primarily determined by the value of C, and vice versa. When rate bits become larger, the overall curve shifts to the lower left, which means we can achieve a better classification and a better distortion level simultaneously.
III-A2 Gaussian source